OpenStudy (anonymous):

Estimate (1.01)6 using differentials (or equivalently, a linear approximation).

OpenStudy (anonymous):

how do i find the f(x)?

OpenStudy (anonymous):

Taylor polynomial?

OpenStudy (anonymous):

what how do you do that?

OpenStudy (anonymous):

A first order Taylor polynomial would give you a linear approximation. Is that supposed to be raised to the power of 6?

OpenStudy (phi):

how about f(x)= x^6 use a=1 (it is close to 1.01, but a^6 is easy to find)